# Decoupled Multimodal Fusion for User Interest Modeling in Click-Through Rate Prediction

### v1

@@ 4.4. Complementary Modality Modeling
Table 1. Statistics of Datasets
Dataset #Users #Items #Samples
Amazon(Electronics) 192k 63k 1.7M
Industry 8.7M 20.9M 469M
5. Offline Experiments
5.1. Datasets
To verify the effectiveness of DMF, we conduct experiments on both public and real-world industrial datasets. For the public dataset, we choose the Amazon [^2] http://jmcauley.ucsd.edu/data/amazon/ dataset, which has been widely used in CTR prediction (Pi et al., 2019; Zhou et al., 2018; Zhang et al., 2025). For the industrial dataset, we use real-world data collected from the Lazada [^3] http://www.lazada.com/ recommendation system for experiments.
The Amazon dataset consists of product reviews and metadata from Amazon. We use the Electronics subset to evaluate DMF. As in (Pi et al., 2019), we treat product reviews as user interaction sequences and organize them chronologically for each user. For a user with a sequence of $T$ interactions, we use the $T$-th interaction as the positive instance and randomly sample another interaction as the negative instance. We also apply 5-core filtering to both users and items.
Basic statistics of these datasets are given in Table 1.
5.2. Evaluation Metrics
For offline experiments, we follow previous work (Zhou et al., 2018; Chang et al., 2023) to adopt the widely used Area Under Curve (AUC) and Group Area Under Curve (GAUC) for evaluation. GAUC is a variant of AUC, which measures the goodness of intra-user order by averaging AUC over users and is shown to be more relevant to online performance. Note that AUC and GAUC reported for Amazon dataset are exactly the same, as this dataset is organized by user group. Therefore, we simply report AUC for Amazon dataset.
Table 2. Model Comparison on Amazon Dataset and Industrial Dataset. We report the mean and standard deviation (std) over 5 runs. $\Delta_{\text{AUC}}$ is calculated to indicate the average performance boost compared with the baseline (SASRec) over datasets. The best and second-best results are marked in bold and underline, respectively. For the industrial dataset, an improvement of 0.1% in AUC and GAUC in the offline evaluation is considered significant for CTR prediction, as observed in (Wang et al., 2021; Chang et al., 2023).
Model Amazon(Electronics) Industry
@@ 5.2. Evaluation Metrics
DMF+MAKE 0.8299 $\pm$ 0.00262 +5.23% 0.6678 $\pm$ 0.00060 +1.87% 0.6195 $\pm$ 0.00082 +1.47%
5.3. Baselines
@@
  • SASRec. A target-agnostic user interest modeling approach that adopts the unidirectional self-attention to capture sequential user interests (Kang and McAuley, 2018).
@@
  • DIN. A target-aware user interest modeling approach that uses activation units to model sequential user interests (Zhou et al., 2018).
@@
  • TA. A target-aware user interest modeling approach derived from standard multi-head attention (Vaswani et al., 2017), which was originally introduced as MHTA in Section 3.2 and is henceforth abbreviated as TA for brevity.
@@
  • $\text{BFS}_{MF}$. An extension of the target attention in TWIN (Chang et al., 2023). TWIN builds a novel attention mechanism via Behavior Feature Splitting (BFS) to efficiently integrate user-item cross features into TA. We adopt the same strategy to incorporate multimodal similarity scores.
@@
  • SIMTIER. A histogram-based approach that converts high-dimensional multimodal representations into a compact similarity histogram. The resulting histogram is treated as a low-dimensional vector and fed into a downstream MLP alongside other embeddings (Sheng et al., 2024).
@@
  • MAKE. A three-stage framework that employs secondary pre-training to enable convergence of multimodal-related parameters, followed by knowledge transfer to downstream tasks (Sheng et al., 2024).
@@
In summary, SASRec, DIN, and TA are ID-based approaches that rely solely on sparse ID features to capture behavioral patterns. In contrast, $\text{BFS}_{MF}$, SIMTIER, and MAKE [^4] Note that DMF is orthogonal to pre-training paradigms. Here we combine DMF with MAKE’s pre-training strategy to show that architectural and training improvements can be combined for further gains. are multimodal approaches that integrate multimodal representations to enhance the effectiveness of user interest modeling.
5.4. Implementation Details
In all experiments, models are trained using the Adam optimizer with an initial learning rate of 0.001 and exponential decay (decay rate: 0.9). ID feature embeddings are set to 32 dimensions. For the Amazon dataset, the attention module uses 4 heads and a hidden size of 32, followed by a fully connected network with dimensions 80 × 40 × 1. The batch size is 128. In the industrial dataset, 128-dimensional visual and textual embeddings are fused into a single 128-dimensional vector for multimodal similarity computation. Although this leads to a slight drop in CTR performance, it significantly reduces storage and computational costs. Here, the attention network has 4 heads and a hidden size of 128, with an MLP of 128 × 64 × 1. The batch size is 1024, and the secondary pre-training stage of MAKE is conducted for 10 epochs.
5.5. Performance Comparison
5.5.1. Overall Performance
The empirical results in Table 2 highlight key insights into user interest modeling with ID-only and multimodal features. Target-agnostic models such as SASRec underperform due to their static user representations that ignore candidate item semantics. In contrast, target-aware models like DIN and TA dynamically reweight historical behaviors based on relevance to the candidate, achieving consistent gains.
Incorporating multimodal signals brings further improvements. Methods including $\text{BFS}_{MF}$, SIMTIER, and MAKE outperform ID-only baselines. These approaches mitigate modality misalignment through similarity encoding or multi-stage pre-training but process ID and multimodal features separately, limiting fine-grained interaction between content semantics and behavioral signals.
Despite DMF+MAKE, DMF achieves the highest performance among all models, demonstrating the effectiveness of its design. DTA, which fuses ID and multimodal signals through decoupled target-aware attention, already outperforms existing methods, highlighting the value of behavior-level modality interaction. The further gain of DMF over DTA confirms that integrating semantic generalization via histogram modeling with fine-grained behavioral modeling strengthens user representation learning.
DMF is complementary to existing training paradigms. When combined with MAKE’s multi-stage pre-training, DMF+MAKE yields further improvements. This shows that structural innovation in fusion architecture and optimization strategies in training can be jointly exploited for better performance.
5.5.2. Ablation Study
@@
Table 3. Ablation Study on Side Information Fusion Strategies. $\Delta_{\text{AUC}}$ is calculated to indicate the average performance boost compared with the baseline (DTA) over the industrial dataset. $\mathcal{C}_{\text{inf}}$ denotes the online inference complexity.
Model AUC (mean $\pm$ std) $\Delta_{\text{AUC}}\uparrow$ GAUC (mean $\pm$ std) $\Delta_{\text{GAUC}}\uparrow$ $\mathcal{C}_{\text{inf}}$
TAearly 0.6644 $\pm$ 0.00212 -0.01% 0.6159 $\pm$ 0.00091 +0.01% $\text{O}(BLd^{2})$
TAlate 0.6615 $\pm$ 0.00101 -0.30% 0.6145 $\pm$ 0.00052 -0.13% $\text{O}(Ld^{2}+BLd)$
DTAnon-invasived 0.6624 $\pm$ 0.00079 -0.21% 0.6129 $\pm$ 0.00063 -0.29% $\text{O}(Ld^{2}+BLd)$
DTA 0.6645 $\pm$ 0.00035 - 0.6158 $\pm$ 0.00043 - $\text{O}(Ld^{2}+BLd)$
We conduct an ablation study to analyze the impact of different side information fusion strategies on both model effectiveness and inference efficiency, as shown in Table 3. The results highlight a clear trade-off between performance and computational cost. $\mathcal{C}_{\text{inf}}$ denotes the dominant term in FLOPs for computing attention. Specifically, TAearly represents early fusion by concatenating ID embeddings and multimodal similarity scores before computing $K$ and $V$. It achieves performance close to our full model (only $0.01\%$ in AUC), demonstrating the potential of fine-grained interaction modeling. However, it suffers from high inference complexity $\mathcal{O}(BLd^{2})$ due to recomputing $K$ and $V$ for each candidate, making it impractical for large-scale deployment.
In contrast, TAlate adopts late fusion, where multimodal features are combined only at the final prediction layer. While computationally efficient ($\mathcal{O}(Ld^{2}+BLd)$), it underperforms DTA by $0.30\%$ in AUC and $0.13\%$ in GAUC, confirming that delaying fusion limits the model’s ability to capture behavioral-level semantic interactions.
To validate the design of our decoupled fusion, we compare against DTAnon-invasived, a variant that does not inject multimodal features into the value ($V$) path—only updating the key ($K$). This version degrades performance by $0.21\%$ in AUC and $0.29\%$ in GAUC, indicating that enriching both $K$ and $V$ with modality-aware signals is crucial for effective interest modeling.
In contrast, our full DTA model achieves the best performance while maintaining low inference complexity $\mathcal{O}(Ld^{2}+BLd)$, thanks to the reuse of ID-derived $K$ and $V$ as target-agnostic nodes. These results demonstrate that DTA successfully balances expressiveness and efficiency through decoupled fusion.
5.5.3. Hyper-parameters Study
@@
Figure 3. Performance with varying representation aggregating hyperparameter $\alpha$. When $\alpha=0$, only the modality-centric modeling strategy is employed, and when $\alpha=1$, only the modality-enriched modeling strategy is utilized.:
@@
The representation aggregating hyperparameter $\alpha$ controls the trade-off between the modality-enriched representation $\text{R}_{me}$ and the modality-centric representation $\text{R}_{mc}$, effectively adjusting the relative contribution of each modeling strategy to the final user interest representation $\text{R}_{u}$. Figure 3 illustrates the sensitivity of DMF to $\alpha$ on Amazon and industrial datasets. We find that the worst performance consistently occurs when $\alpha=0$, as in this setting, DMF relies solely on $\text{R}_{mc}$, completely disregarding the fine-grained behavioral signals captured by $\text{R}_{me}$, indicating that content-driven generalization alone, while robust, is insufficient for optimal personalization. As $\alpha$ increases, the influence of $\text{R}_{me}$ grows, leading to consistent improvements in prediction performance. The Amazon dataset shows the peak performance at $\alpha=0.3$, and the industrial dataset shows the peak performance at $\alpha=0.7$, suggesting that the optimal balance favors $\text{R}_{me}$ but still requires substantial contribution from $\text{R}_{mc}$ to maintain semantic coherence and generalization capability. At $\alpha=1$, DMF depends entirely on $\text{R}_{me}$, which fuses item identity signals with multimodal context. However, ID-based embeddings embedded within this modality-enriched modeling strategy can be biased or stale, especially under interest shift or adversarial behavior, leading to suboptimal performance.
6. Online Experiments
@@
We apply the DMF framework to the ranking model in the product recommendation system of the international e-commerce platform Lazada, ventures of Thailand. Each candidate product in the ranking system will be ranked based on a rank score, which is estimated based on the CTR and CVR (Conversion Rate).
@@
The online base model used TA to extract user interests from user interaction sequences and used SIMTIER to integrate the pre-trained multimodal representations, including the representations for both visual and textual modalities. Compared with the base model, we substitute the TA and SIMTIER modules with DMF to model behavioral patterns and multimodal signals together. After conducting an online A/B test for 12 days, DMF demonstrated promising results, achieving relative improvements of 5.30% in CTCVR and 7.43% in GMV. The deployment introduces moderate additional computation per request; however, this is efficiently absorbed through horizontal scaling without increasing end-to-end latency. The associated increase in resource cost is marginal compared to the magnitude of business gains, demonstrating DMF’s favorable return on investment and deployability at scale.
7. Conclusion
@@
We proposed Decoupled Multimodal Fusion (DMF), a hybrid framework for user interest modeling in CTR prediction. By introducing Decoupled Target Attention (DTA), DMF enables fine-grained interaction between ID-based and multimodal representations while preserving inference efficiency through computational decoupling. The Complementary Modality Modeling (CMM) module fuses modality-centric and modality-enriched representations for more comprehensive interest modeling. Experiments show that DMF outperforms state-of-the-art methods on both public and industrial datasets. Deployed on Lazada’s recommendation system, DMF achieves significant improvements of +5.30% in CTCVR and +7.43% in GMV, demonstrating its effectiveness and practical value in real-world applications.

### v2

@@ IV-E Discussion
V Experiments
In this section, we conduct extensive experiments to answer the following research questions:
  • RQ1: How does the proposed DMF framework perform compared with the state-of-the-art CTR prediction baselines?
  • RQ2: How do different side information fusion strategies affect model performance and inference efficiency?
  • RQ3: How sensitive is DMF to the representation aggregation hyperparameter $\alpha$?
  • RQ4: How does the interaction sequence length influence the optimal trade-off between modality-centric and modality-enriched modeling?
  • RQ5: Can DMF capture meaningful user interest patterns in real-world scenarios?
  • RQ6: How does the proposed DMF framework perform on the real-world e-commercial platform?
TABLE II: Model Comparison on Amazon Dataset and Industrial Dataset. We report the mean and standard deviation (std) over 5 runs. $\Delta_{\text{AUC}}$ is calculated to indicate the average performance boost compared with the baseline (SASRec) over datasets. The best and second-best results are marked in bold and underline, respectively. For the industrial dataset, an improvement of 0.1% in AUC and GAUC in the offline evaluation is considered significant for CTR prediction, as observed in [38, 13].
Model Amazon Industry
@@ V Experiments
V-A Datasets
To verify the effectiveness of DMF, we conduct experiments on both public and real-world industrial datasets. For the public dataset, we choose the Amazon∗*∗**http://jmcauley.ucsd.edu/data/amazon/ dataset, which has been widely used in CTR prediction [39, 5, 11]. For the industrial dataset, we use real-world data collected from the Lazada∗*∗**http://www.lazada.com/ recommendation system for experiments.
The Amazon dataset consists of product reviews and metadata from Amazon. We use the Electronics subset to evaluate DMF. As in [39], we treat product reviews as user interaction sequences and organize them chronologically for each user. For a user with a sequence of $T$ interactions, we use the $T$-th interaction as the positive instance and randomly sample another interaction as the negative instance. We also apply 5-core filtering to both users and items.
Basic statistics of these datasets are given in Table III.
TABLE III: Statistics of Datasets
Dataset #Users #Items #Samples
Amazon 192k 63k 1.7M
Industry 8.7M 20.9M 469M
V-B Evaluation Metrics
For offline experiments, we follow previous work [5, 13] to adopt the widely used Area Under Curve (AUC) and Group Area Under Curve (GAUC) for evaluation. GAUC is a variant of AUC, which measures the goodness of intra-user order by averaging AUC over users and is shown to be more relevant to online performance. It is worth mentioning that AUC and GAUC reported for the Amazon dataset are the same, as this dataset is organized by user group.
V-C Baselines
@@
  • SASRec. A target-agnostic user interest modeling approach that adopts the unidirectional self-attention to capture sequential user interests [19].
@@
  • DIN. A target-aware user interest modeling approach that uses activation units to model sequential user interests [5].
@@
  • TA. A target-aware user interest modeling approach derived from standard multi-head attention [32], which was originally introduced as MHTA in Section III-B and is henceforth abbreviated as TA for brevity.
@@
  • $\text{BFS}_{\text{MF}}$. An extension of the target attention in TWIN [13]. TWIN builds a novel attention mechanism via Behavior Feature Splitting (BFS) to efficiently integrate user-item cross features into TA. We adopt the same strategy to incorporate multimodal similarity scores.
@@
  • SIMTIER. A similarity histogram-based approach that converts high-dimensional multimodal representations into a compact similarity histogram. The resulting histogram is treated as a low-dimensional vector and fed into a downstream MLP alongside other embeddings [9].
@@
  • MAKE. A three-stage framework that employs secondary pre-training to enable convergence of multimodal-related parameters, followed by knowledge transfer to downstream tasks [9].
@@
In summary, SASRec, DIN, and TA are ID-based approaches that rely solely on sparse ID-based features to capture behavioral patterns. In contrast, $\text{BFS}_{\text{MF}}$, SIMTIER, and MAKE are multimodal approaches that integrate multimodal representations to enhance the effectiveness of user interest modeling.
@@
TABLE IV: Ablation Study on Side Information Fusion Strategies. $\Delta_{\text{AUC}}$ is calculated to indicate the average performance boost compared with the baseline (DTA) over the industrial dataset.
Model AUC (mean $\pm$ std) $\Delta_{\text{AUC}}\uparrow$ GAUC (mean $\pm$ std) $\Delta_{\text{GAUC}}\uparrow$
[…]
DTA 0.6645 $\pm$ 0.00035 - 0.6158 $\pm$ 0.00043 -
V-D Implementation Details
In all experiments, models are trained using the Adam optimizer. ID feature embeddings are set to 32 dimensions. For the Amazon dataset, the attention module uses 4 heads and a hidden size of 32, followed by a fully connected network with dimensions 80 × 40 × 1. The batch size is 128. In the industrial dataset, 128-dimensional visual and textual embeddings are fused into a single 128-dimensional vector for multimodal similarity computation. Although this leads to a slight drop in CTR performance, it significantly reduces storage and computational costs. Here, the attention network has 4 heads and a hidden size of 128, with an MLP of 128 × 64 × 1. The batch size is 1024, and the number of buckets $M$ is 35. The secondary pre-training stage of MAKE is conducted for 10 epochs.
V-E Performance Comparison
V-E1 Overall Performance(RQ1)
The empirical results in Table II highlight key insights into user interest modeling with ID-only and multimodal features. Target-agnostic models such as SASRec underperform due to their static user representations that ignore target item semantics. In contrast, target-aware models like DIN and TA dynamically reweight historical interactions based on relevance to the target item, achieving consistent gains.
Incorporating multimodal signals brings further improvements. Methods including $\text{BFS}_{\text{MF}}$, SIMTIER, and MAKE outperform ID-only baselines. These approaches mitigate modality misalignment through similarity encoding or multi-stage pre-training but process ID and multimodal features separately, limiting fine-grained interaction between content semantics and behavioral signals.
DMF achieves the highest performance among all models, demonstrating the effectiveness of its design. DTA, which fuses ID and multimodal signals through decoupled target-aware attention, already outperforms existing methods, highlighting the value of behavior-level modality interaction. The further gain of DMF over DTA confirms that integrating semantic generalization via histogram modeling with fine-grained behavioral modeling strengthens user representation learning.
V-E2 Ablation Study(RQ2)
[…]
Among all variants, DTA attains the best AUC and competitive GAUC, while maintaining significantly lower inference latency. This balance of accuracy and efficiency makes DTA particularly suitable for real-world deployment. To evaluate inference efficiency under real-world serving conditions, we perform stress tests on a single NVIDIA A10 GPU with a hidden dimension of 128, user sequence length of 400, and 1200 candidate target items per request, which is representative of industrial-scale ranking workloads. In this setting, the decoupled fusion strategy achieves a 3× higher QPS compared to early fusion, highlighting its strong suitability for high-scale, latency-sensitive recommendation systems.
@@ V-E2 Ablation Study(RQ2)
Figure 5: Performance with varying representation aggregating hyperparameter $\alpha$. When $\alpha=0$, only the modality-centric modeling strategy is employed, and when $\alpha=1$, only the modality-enriched modeling strategy is utilized.
V-E3 Hyper-parameters Study(RQ3)
@@
The representation aggregating hyperparameter $\alpha$ controls the trade-off between the modality-enriched representation $\text{R}_{me}$ and the modality-centric representation $\text{R}_{mc}$, effectively adjusting the relative contribution of each modeling strategy to the final user interest representation $\text{R}_{u}$. Fig. 5 illustrates the sensitivity of DMF to $\alpha$ on Amazon and industrial datasets. We find that the worst performance consistently occurs when $\alpha=0$, as in this setting, DMF relies solely on $\text{R}_{mc}$, completely disregarding the fine-grained behavioral signals captured by $\text{R}_{me}$, indicating that content-driven generalization alone, while robust, is insufficient for optimal personalization. As $\alpha$ increases, the influence of $\text{R}_{me}$ grows, leading to consistent improvements in prediction performance. The Amazon dataset shows the peak performance at $\alpha=0.3$, and the industrial dataset shows the peak performance at $\alpha=0.7$, suggesting that the optimal balance favors $\text{R}_{me}$ but still requires substantial contribution from $\text{R}_{mc}$ to maintain semantic coherence and generalization capability. At $\alpha=1$, DMF depends entirely on $\text{R}_{me}$, which fuses item identity signals with multimodal context. However, ID-based embeddings embedded within this modality-enriched modeling strategy can be biased or stale, especially under interest shift or adversarial behavior, leading to suboptimal performance.
Figure 6: Relationship between user interaction sequence length and model performance. (a) Proportion of samples for users with short ($L\leq 10$) and long ($L>10$) interaction sequences on Amazon and Industry datasets.(b) GAUC improvements achieved by DMF over DTA for users with short ($L\leq 10$) and long ($L>10$) interaction sequences on Amazon and Industry datasets.
Figure 7: Illustration of the attention weights in TA and DTA. Interactions with high relevance to the target item get high attention weights.
V-E4 Analysis of Representation Aggregation Dynamics(RQ4)
An interesting question is why the optimal value of $\alpha$ varies significantly between Amazon and industrial datasets. Fig. 6 reveals that the modality-centric representation $\mathbf{r}_{\mathrm{mc}}$ delivers substantially higher GAUC gains for users with short interaction sequences ($L\leq 10$) than for those with long sequences ($L>10$), consistently across both datasets. This effect explains the divergent optimal $\alpha$ values: the Amazon dataset, where 84.75% of users are low-activity ($L\leq 10$), achieves peak performance at $\alpha=0.3$, while the industrial dataset, dominated by high-activity users (89.35% with $L>10$), favors $\alpha=0.7$ to prioritize modality-enriched modeling.
This pattern suggests that semantic generalization is more effective under conditions of action sparsity. For low-activity users, limited historical interactions constrain the reliability of behavior-driven personalization; here, modality-centric features provide a robust semantic prior that mitigates data scarcity. In contrast, high-activity users generate abundant behavioral signals, enabling modality-enriched representations to precisely align recommendations with fine-grained preferences by fusing ID-based collaboration with multimodal context. Thus, the optimal fusion weight $\alpha$ is not an arbitrary hyperparameter but a reflection of the user population’s activity profile.
From a system design perspective, this insight advocates for activity-aware fusion strategies in industrial recommenders. Rather than using a global $\alpha$, platforms should adapt the aggregation weight based on user activity level, that is, assigning a higher weight to $\mathbf{r}_{\mathrm{mc}}$ for new users or low-activity user groups and to $\mathbf{r}_{\mathrm{me}}$ for high-activity user groups. Such an approach would dynamically balance robustness and personalization according to the informational sufficiency of each user interaction history, transforming a static model component into an intelligent, context-responsive mechanism that serves diverse user groups effectively.
V-E5 Case Study(RQ5)
We conduct case study to reveal the inner structure of DTA on the industrial dataset. Since attention weights computed via softmax normalization are not comparable across models, we use sigmoid normalization to compute attention weights for both TA and DTA. Fig. 7 visualizes the attention weights assigned by TA and DTA to user interaction traces with respect to a target item.
As shown in Fig. 7, TA assigns the highest weight $0.2965$ to Cotton Fabric, which is a highly popular item with frequent interactions indicated by three flame icons. However, this item exhibits low semantic relevance to the target item, which is an assembled teak cabinet. In contrast, Teak Display Cabinet, a less frequently interacted item indicated by one flame icon, shares high multimodal similarity $0.9243$ with the target item, yet receives only $0.2127$ attention weight under TA, reflecting its inability to capture content-level relevance.
DTA integrates multimodal cosine similarity as a target-aware signal to modulate attention computation. DTA correctly identifies that Teak Display Cabinet is semantically aligned with the target item and assigns it the highest attention weight ($0.6586$), significantly surpassing all other items in the sequence. This demonstrates the capability of DTA to transcend popularity bias and prioritize behaviorally rare but semantically relevant interactions.
V-E6 Online A/B Testing (RQ6)
@@
We apply the DMF framework to the ranking model in the product recommendation system of the international e-commerce platform Lazada, ventures of Thailand. Each target item in the ranking system will be ranked based on a rank score, which is estimated based on the CTR and CVR (Conversion Rate).
@@
The online base model used TA to extract user interests from user interaction sequences and used SIMTIER to integrate the pre-trained multimodal representations, including the representations for both visual and textual modalities. Compared with the base model, we substitute the TA and SIMTIER modules with DMF to model behavioral patterns and multimodal signals together. After conducting an online A/B test for 12 days, DMF demonstrated promising results, achieving relative improvements of 5.30% in CTCVR and 7.43% in GMV. The deployment introduces moderate additional computation per request; however, this is efficiently absorbed through horizontal scaling without increasing end-to-end latency. The associated increase in resource cost is marginal compared to the magnitude of business gains, demonstrating the favorable return on investment and deployability at scale of DMF.
TABLE V: Online A/B testing results. The improvements on GMV and CTCVR are reported.
Metric GMV CTCVR
Lift rate +7.43% +5.30%
VI Conclusion
@@
We proposed Decoupled Multimodal Fusion (DMF), a hybrid framework for user interest modeling in CTR prediction. By introducing Decoupled Target Attention (DTA), DMF enables fine-grained interaction between ID-based and multimodal representations while preserving inference efficiency through computational decoupling. The Complementary Modality Modeling (CMM) module fuses modality-centric and modality-enriched representations for more comprehensive interest modeling. Experiments show that DMF outperforms state-of-the-art methods on both public and industrial datasets. Deployed on the product recommendation system of the international e-commerce platform Lazada, DMF achieves significant improvements of +5.30% in CTCVR and +7.43% in GMV, demonstrating its effectiveness and practical value in real-world applications.